OpenStudy (anonymous):

find the x value for the point that splits segment CD in half if point C is located at (-2, 4) and point D is located at (3, 7)

OpenStudy (johnweldon1993):

So the midpoint would be found using the formula \[\large \frac{x_1 + x_2}{2}\] So you take your 2 'x' coordinate points...(-2 and 3) add them, and then divide the answer by 2 \[\large \frac{-2 + 3}{2} = \frac{1}{2}\] So 1/2 or .5 would be the 'x' value you are looking for